OpenStudy (lovelycharm):

Producing a musical cost $88,000 $5900 per perfomance. One sold-out perfomaance earn $7500 in revenue. If every perfomance sell out, how many performance are needed to break even?

OpenStudy (anonymous):

x stands for # of performances 7500x = 88,000 solve for x... can you take it from here ?

OpenStudy (lovelycharm):

@kelliegirl33 so u add or subtract to find x

OpenStudy (anonymous):

you divide by 7500 on both sides.

OpenStudy (lovelycharm):

so it 80500

OpenStudy (anonymous):

One thing you need to learn is that when the problem is being multiplied, like this one....(7500)(x), you divide. If the problem was 7500 + x = 88,000, then you can subtract....just remember when it is multiplied, you divide. Your answer is incorrect....let me show you... 7500x = 88,000 7500x/7500 = 88,000/7500 (the 7500 on the right side, cancels out) x = 11.7 rounded = 12 12 performances are needed to break even.

OpenStudy (anonymous):

oops....the 7500 on the left side cancels out

OpenStudy (lovelycharm):

okay

OpenStudy (anonymous):

Do you understand ?

OpenStudy (lovelycharm):

not really

OpenStudy (anonymous):

7500 is what is earned for a sold out performance. How many sold out performance will equal 88,000.? I set up the equation 7500x = 88,000 where x is the number of performances. Then you have to solve for x by getting x on one side and everything else on the other side. To get rid of the 7500 on the left side we have to divide by 7500 on both sides. This will make the 7500 on the left side cancel out. But what you do on one side, you have to do on the other side. So by dividing 88,000 by 7500, you get 11.7. There cannot be 11.7 performances so I rounded up to the next whole number which is 12......12 x 7500 = 90,000. Eleven performances will not work because 11 x 7500 = 82,500 and that is not enough to break even.

OpenStudy (lovelycharm):

what the answer is so it 12

OpenStudy (anonymous):

yes...the answer is 12

OpenStudy (lovelycharm):

thank u for caring and helping

OpenStudy (anonymous):

no problem....glad I could help :)